OpenStudy (anonymous):

Let P(t) be the population of a country, in millions, t years after 1990, with P(6) = 3.21 and P(11) = 3.69. Find a formula for P(t) assuming that it is exponential.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

What type of class is this for? :) I don't want to start with it in differential equation form if you're taking algebra or something similar ^^ heh

OpenStudy (anonymous):

Its algebra. Im pretty sure im supposed to use \[y=c*b^x\]

OpenStudy (anonymous):

I have to figure out c and b with the two points given

zepdrix (zepdrix):

Ok let's use the information given to FIRST find b. It's going to be a little bit confusing since we'll do a little bit more than simply add or subtract equations. :D Let's first start with our equation like this, since it's in terms of Population and time.\[\huge P(t)=Cb^t\]

zepdrix (zepdrix):

\[\large P(6)=3.21 \qquad \rightarrow \qquad 3.21=Cb^{6}\]\[\large P(11)=3.69 \qquad \rightarrow \qquad 3.69=Cb^{11}\]Understand how I got these two equations? Plugging 6 and 11 into our model.

OpenStudy (anonymous):

Affirmative

zepdrix (zepdrix):

From here we can DIVIDE these equations to ELIMINATE the C term. Which will allow us to solve for b.\[\large \frac{P(6)}{P(11)} \qquad \rightarrow \qquad \frac{3.21}{3.69}=\frac{Cb^6}{Cb^{11}}\]I probably should have divided them the other way, with the 11 on top :) But oh well, it'll still work out.

OpenStudy (anonymous):

I didnt know you could do that =p so it gives 3.21/3.69=b^-5?

zepdrix (zepdrix):

yes, from there let's change it to a +5 then take the 5th root of both sides.

zepdrix (zepdrix):

I don't just mean, "change it", I mean find the equation in terms of b^5, heh. If that part is confusing, we can do the steps :D

OpenStudy (anonymous):

Hmmm im not sure...would i change it to like 1/b^5?

zepdrix (zepdrix):

Yes, then you can rewrite BOTH sides as reciprocals :)

zepdrix (zepdrix):

If that is confusing, you can always just cross multiply or something similar ^^

OpenStudy (anonymous):

yeah not sure what ya mean there. 3.69=3.21b^5 3.69/3.21=b^5 \[\sqrt[5]{3.69/3.21}\] did i do that right?

zepdrix (zepdrix):

Yah looks good :) Now the only problem we'll have from here is.... If we want to accurately get C, we'll want to leave b like that for now. We don't want to round at this point, since we still need to solve for another variable.

zepdrix (zepdrix):

From this point, it shouldn't be too bad. Let's plug our b value into P(6) or P(11) and solve for C.

zepdrix (zepdrix):

C represents the INITIAL POPULATION. So hopefully we'll get a number that makes sense :o

OpenStudy (anonymous):

Hopefully. \[3.21=c*\sqrt[5]{3.69/3.21}^{6}\]

OpenStudy (anonymous):

that looks funky

zepdrix (zepdrix):

\[\large P(6)=3.21 \qquad \rightarrow \qquad 3.21=C\left(\sqrt[5]{\frac{3.69}{3.21}}\right)^6\]Hehe it sure does :3 you can type "large" or "huge" (without the quotes) in your equation tool to make it a little larger if you need :D cause the exponents can be difficult to read sometimes.

OpenStudy (anonymous):

Didnt know that either =p Do you subtract the fifth root and the exponent 6?

zepdrix (zepdrix):

Hmm this is giving us a really bad number :( Are you sure we used the correct model? I'm more familiar with seeing this model for population.\[\huge P(t)=Ce^{kt}\]But if that's not one that you've been using, then ignore it.

OpenStudy (anonymous):

oopsy. :(

OpenStudy (anonymous):

How do i know which one to use? i assumed it was the first one because thats what we learned in class today, and this problem is on the homework

OpenStudy (anonymous):

But we used the one you just typed a few days ago

OpenStudy (anonymous):

And i thought c was initial population but the problem doesnt say the initial population

OpenStudy (anonymous):

Wait...i used decimals for b and c and typed the final function into my homework and it said it was correct. So we did it right :) well mostly you lol. Thank you

zepdrix (zepdrix):

Oh weird! :) Yes C should represent our initial population. But it's giving me a value of like 2.71... So the oh oh oh oh... it's in MILLIONS. So the initial population was 2.7 million. Ok that makes A LOT more sense! :) I guess we were doing it correctly. I'm sorry about the slow response. I was trying to work on multiple problems at once, and the site is lagging reallyyy bad right now :( hard to move back and forth.

OpenStudy (anonymous):

It's okay. Makes tons more sense. Thanks again =D